HUBBLE SPACE TELESCOPE DAILY REPORT #2043 PERIOD COVERED: 0000Z (UTC) 01/20/98 - 0000Z (UTC) 01/21/98 Daily Status Report as of 021/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7900 (STIS CCD Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, observations with the NIC/3 Camera. This proposal will run during the orbits devoted to NICMOS Hubble Deep Field (HDF) prime proposal 7235. The observations were completed as scheduled, and no anomalies were reported. 1.3 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of STIS/MA1/MA2 7921 (STIS MAMA Pure Parallels for NIC/3 HDF - Spec) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make pure parallel, ultraviolet observations to be used in the NIC/3 HDF campaign. The MAMA images will be used to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.6 Completed Two Sets of STIS/MA1/MA2 7920 (STIS MAMA Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (MA1 AND MA2) was used to make parallel field, ultraviolet imaging to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were completed as planned, and no problems were reported. 1.7 Completed NIC/3 7815 (NICMOS Earth Flats) The Near Infrared Camera (Camera Three) was used to make a Earth flat field calibration image. This proposal obtains a series of earth flats for a subset of narrow and medium band NIC/3 filters in advance of the NIC/3 campaign. These flats will be used to generate the master flats for narrow-band filters and to determine the OTA illumination pattern as a function of wavelength. The observations were completed as planned, and no anomalies were reported. 1.8 Completed NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/1/2/3 7223 (Spectrophotometry and Imaging of Pluto and Charon) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of Charon. Detailed spectra of the individual members of the Pluto-Charon system are not known. The observations were completed as scheduled, and no problems were reported. 1.10 Completed NIC/1 7693 (NICMOS Pupil Transfer Function) The Near Infrared Camera (Camera One) was used to make observations of P177D in order to perform pupil transfer function observations. The output of this program will be a two-dimensional surface fit that will give the large scale corrections needed to flat-field the images to an accuracy of about 2%. Also, changes in the PSF will be measured as a function of position within the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxy NGC-6207 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. As described in 2.1, this proposal executed on gy control only. Otherwise, the observations were executed as planned, and no further anomalies were reported. 1.12 Completed Two Sets of NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the protoplanetary disk around star CI-Tau to search for water in their disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed NIC/1/2/3 7239 (Surface Compositions of Kuiper Belt Objects) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Kuiper Belt object 1996TL66 in order to study its surface composition. The surfaces of Kuiper Belt Objects [KBO's] contain a record of formative and dynamical processes in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 For the acquisition at 020/2030Z, both FGS-1 and FGS-2 acquired fine lock a total of five times, each time dropping lock a short time later. Finally, the proposal described in 1.11 executed on gyro control only. HSTAR 6324 was written. Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: It was incorrectly reported in yesterday's report that the STIS MAMA2 reset occurred while the high voltage was off. In fact, the status buffer message at 020/080058Z indicated that the high voltage had anomalously been turned off. STIS MAMA2 recovery was accomplished via ROP NS-14 at 020/1718Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH